The area of a rectangle is 168 square ft. Determine the length, width, and diagonal of a rectangle if the length is 3 feet greater than 3 times the width.
Answer by ewatrrr(24785)


Hi,
Let x represent the width. then the length is (3x+3)
Area is l*w
(3x+3)*x == 168ft^2
3x^2 + 3x = 168ft2
3x^ + 3x - 168 = 0
x^ + x - 56 =0
factoring
(x - 8)(x+7)= 0
x = 8 the width is 8ft
x = -7 cannot use
width is 8 and the length is 27 (3*8 + 3)
finding d the diagonal implementing the Pythagorean Therem
8^2+ 27^2 = d^2
d = 28.160 ft